# Customizing Motion in Text-to-Video Diffusion Models

### v1

@@ 3.2 Approach for Motion Customization
| Real (Ours) | 70.6 | 8.7 |
4 Experiments
In this section, we quantitatively evaluate our approach (Section 4.1), perform an ablation study (Section 4.2), and show qualitative results (Section 4.3).
Implementation details. In all our experiments we use the ZeroScope text-to-video diffusion model [3] as the pre-trained network. Our implementation is building on the public repository [8]. In our concept exemplar dataset , we apply spatial and color augmentation on the video examples. We train the model with a mini-batch size 1 on a single A100 Nvidia GPU. We train the models up to 3000 training steps with learning rate .
4.1 Quantitative Evaluation
We quantitatively evaluate our approach by computing metrics corresponding to the quality of the generated motion and the overall fidelity of the generated video with respect to the input text prompt.
Dataset. We quantitatively evaluate our approach on the Jester Dataset [27]. The dataset contains 148,092 crowd-sourced videos of 27 human gestures by diverse actors performing a gesture in front of a static camera. We select 5 gesture classes as our target motions – Swiping Down, Drumming Finders, Rolling Hand Forward, Shaking Hand, Sliding Two Fingers Up. From each class, we choose at random 10 videos from the validation set as our training set. To learn the motion, we caption the videos with a generic text description “a person doing a V* sign”. As a regularization set for the motion learning, we also choose 50 videos from each of the dataset classes Doing Other Things, No gesture and assign them text captions “a person doing other things” and “a person in front of a computer”. We design a test set containing 100 text prompts that detail a person’s appearance, for example “A female firefighter doing the V* gesture” or “A Japanese toddler doing the V gesture”. We fix 3 random seeds per text prompt to evaluate our experiments.
Motion accuracy metric. The Jester dataset has been widely adopted for the task of action recognition. We leverage this for our motion accuracy metric, measuring gesture recognition accuracy with a video classifier, pre-trained on the Jester training set [49]. The classifier achieves 94.78% accuracy across all gesture classes, making it suitable for our evaluation. Note that the simple test set prompts typically result in generated videos that adhere to the spatial layout of the gesture recognition classifier’s training videos with a single person, making them in-distribution.
Text alignment metrics. To regulate overfitting, we measure how faithfully the model adheres to the appearance described in the text prompt with an appearance score defined as the mean CLIP score between the generated frames and the part of the text prompt describing the person’s appearance (e.g., “A female firefighter”). We empirically find that the appearance score is a good proxy to determine when the model starts to overfit. For a fair comparison across all models, we fine-tune each model until this appearance score reaches a chosen value (here 0.265), determined empirically based on the visual quality of the results. Please refer to the Supplemental Material for an additional discussion on the tradeoff between motion accuracy and appearance alignment.
To assess how well the generated video corresponds to the overall text prompt (both motion and appearance), we measure the text alignment score as a mean CLIP score between the video frames and a full caption written in natural language. For example, for the class Swiping Down and a text prompt “A female firefighter doing the V* gesture”, a full caption would be “A female firefighter swiping hand down.” We use this score as another criterion to compare different methods.
Comparison with image customization methods. We select three well-established image customization approaches: Textual Inversion [9], Dreambooth [32] and Custom Diffusion [22], and adapt them to the motion customization task. Textual Inversion [9] optimizes a text token embedding to learn a novel subject. For Dreambooth [32], we optimize all spatial layers of the model and use a synthetic video regularization set following their prior preservation loss. Custom Diffusion trains the key and value projection matrices in the cross-attention layers of text-to-image diffusion models, optimizes a text token embedding, and uses a regularization set containing real videos. We report the motion accuracy and text alignment metrics in Table 1 and show qualitative comparison in Figure 3. We observe that Textual Inversion completely fails to learn the motion. We hypothesize that this failure is because the text encoder has been trained only on image-text pairs, and the embedding fails to generalize to the video domain. Additionally, the text embedding is injected only into spatial layers in the cross-attention layers, and not the temporal ones. Alternatively, because we are learning an unseen motion pattern, which is more complex than a static object, the embedding does not have enough parameters to learn it. On the other hand, Dreambooth and Custom Diffusion learn to adhere to the spatial structure of the training videos and some hand manipulation. However they fail to accurately reproduce the motion and produce a temporally smooth video. In contrast, our approach yields over a 2 improvement in motion accuracy and more faithfully adheres to the text prompt.
4.2 Ablation study
We perform an ablation study of our method’s components and report the motion accuracy metric in Table 2. In addition, we report a copying score (“Copy”) that measures how much of the training data appearance leaks into the generated data. We use the SSCD description score [29] to detect image copying between the training videos and the generated videos. We compute our copying score as the maximum SSCD description score between the generated frames of the video and first frames of the training examples, and report the percentage of samples that are above a specified threshold. For our experiments, we manually determine a threshold of 0.25, which corresponds to significant copying of the appearance of the person or the background (please see examples in the Supp. Mat.).
@@ 4.2 Ablation study
First, we study the choice of the optimized parameters. As in the case of the image customization methods, it is clear that optimizing the spatial layers of the model is not enough to learn a novel motion pattern, as shown in Table 1(b). We can further observe that training the key and value projections in the cross-attention layers of the spatial layers achieves a good balance in terms of accuracy and copying score, as shown Table 1(a). We observe a two-fold reduction in the copying score when training only those parameters and an 8% improvement in motion accuracy compared to not training any spatial parameters.
@@
Next, in Table 1(c), we consider the role of optimizing the text token embedding. We notice that when jointly training the text token embedding with the model’s parameters, the motion accuracy is higher than when training only the model parameters. However, it also leads to nearly three times as much memorization of the training video appearance, as evidenced by the significantly increased Copy score. We hypothesize that this indicates that the text token embedding is learning something related to the appearance of the training videos.
Consequently, we test our sampling strategy. We compare a model trained with our motion pattern sampling strategy (“Coarse-noise”) to a model that simply samples diffusion time steps uniformly (“Uniform”). As shown in Table 1(d), our sampling strategy improves the motion accuracy and reduces the appearance copying. Following our intuition regarding learning motion patterns rather than fine-grained details, our model is less prone to replicating elements from the training examples.
Lastly, we study the effect of the regularization prior in Table 1(e). We train three models with the same design choice as our method, yet we choose the regularization set to be either: (i) empty, (ii) a synthetic set of videos generated from the original text-to-video model, or (iii) containing real videos. We observe a significant drop in performance both without using a prior as well as using a synthetic prior. Interestingly, despite all settings using the same fine-tuning set, a high-quality regularization set containing real videos improves the overall training. Overall, our method achieves a good balance between high motion accuracy and low copying score.
4.3 Qualitative experiments
To further demonstrate the effectiveness of our approach, we choose the motions: Dab, Airquotes, and 3D rotation, which the original model does not know or cannot reliably perform. For the first two motions, we curate videos from the internet searching for the description of the motion. We use between 5 and 6 videos for each motion and caption the videos with a generic text description such as “A person doing the V* sign” or “A person V*”. To learn 3D rotation, we select 10 sequences of distinct objects from the CO3D Dataset [30] and caption the videos with the text prompt “A camera V* around an [object]”, where [object] is an object class. The results are shown in Figure 4. For visualization, we query for objects not present in the training videos. For the regularization set, we select 100 video and text pairs from the WebVid-10M Dataset [1]. We fine-tune the models for 2000 steps. Please see our companion webpage for additional results and comparison with the pre-trained model.
Motion and appearance customization. We combine the appearance of a specific person with a gesture motion Drumming Fingers from the Jester Dataset. We use the publicly available images of a particular person from the CustomConcept101 Dataset [22], and annotate them with generic text captions {“A X* man”, “A close-up of a X* man”, “A X* man smiling.”}. For the motion dataset, we choose the same 10 videos as in Section 4.1 and a generic text description “a person doing a V* sign”. Our qualitative results are shown in Fig. 6.
Comparison with a reference-based method. We compare our method with Tune-A-Video [45], a video editing method that, given a source video, inflates a text-to-image model to learn a motion pattern from a single video. This method allows generation of novel videos with the same pattern but varying appearance. We adapt Tune-A-Video to fine-tune from multiple videos using the same training set as our setup. In Figure 5 we compare the two methods. We test the generalization abilities of the two methods, firstly combining the novel motion with a different motion: “doing the V* gesture while eating a burger with other hand”. We can see that our method can generate the two motions. Next, we test the motion variation. We show that we can manipulate the execution of the motion with a natural language description; here we specify “doing the V* gesture very slowly and precisely”. Lastly, we show that our method can generalize to multiple people doing the motion while preserving the temporal consistency of the video. The Tune-A-Video baseline produces qualitatively worse results, as can be seen in Figure 5 and the supplementary video.
@@ 5 Conclusion
We present an approach for motion customization in text-to-video diffusion models. Our method can learn a new motion from a set of few exemplar videos of different subjects performing the same gesture. We conduct a thorough ablation study that identifies the key components of the method and evaluate them with respect to motion accuracy. We demonstrate qualitative results of our method, specifically in scenarios where our method generalizes the motion to unseen actors, multiple people performing the motion, and different viewpoints. Additionally, we show how our method can be extended to multi-modal motion and appearance customization to generate videos with custom subjects performing custom motions.
Supplementary Material
This supplemental contains the following material:
Accompanying webpage of qualitative results of our method, the videos shown in the main paper and comparison with the pre-trained text-to-video generations.
(Section A) We provide additional implementation details regarding our quantitative experiments.
(Section B) We discuss the tradeoff between the motion accuracy, appearance alignment and copying score across training iterations.
(Section C) We show qualitative examples along with their copying score.
(Section D) We show the original model’s (i.e. before customization) performance on the Jester dataset.
Appendix A Implementation Details
Appendix B Tradeoff between motion accuracy, appearance alignment, and copying score
When customizing a model with a new motion, we observe a natural bias-variance tradeoff. The longer we train the model, the better motion accuracy becomes, we memorize elements of the training examples, and the model becomes less diverse. We illustrate this in Figure 7. With more training iterations, the network memorizes the gesture as well as visual elements of the training examples. These findings are indicated by both the motion accuracy and copying score increasing. On the other hand, the CLIP appearance score, decreases. The CLIP appearance is defined as an average CLIP score between the part of the text prompt that describes a person’s appearance and the video frames (i.e. in a prompt “A female firefighter doing the V* sign”, the CLIP appearance score accounts for “A female firefighter”). This score is a proxy for visual diversity, since our test set contains visual description of a person, it measures how accurate the description is. As we increase the number of iterations, the videos become less diverse, hence the score decreases. This trend is persistent across the models from our ablation study. To find a balance between those factors, we perform early stopping for all models when the CLIP appearance score reaches 0.265 (i.e. when the image still reasonably well corresponds to the text description).
Appendix C Copying score
Copying score, as defined in 4.2 measures how much of the training data appearance leaks into the generated data. The score corresponds to the percentage of examples in the test set that are above a specified threshold of the SSCD score with any of the training examples. We find that the SSCD copy detection score [29] is a good proxy for determining the memorization of the training examples’ appearance. We measure the SSCD detection as a maximum score between any frame of the training videos and any video frame from the test set. We empirically determine that pairs of videos with a score greater than 0.25 share significant visual similarities. As seen in Figure 8, the similarities can be either the background or the appearance of the person performing the gesture.
Appendix D Pre-trained model’s performance on motions from the Jester dataset.
To test whether the gestures from the Jester dataset, are well-represented in the pre-trained text-to-video model, we evaluate the model with a similar experiment as our ablation experiment (4.2). We query the pre-trained text-to-video model with the prompts from the test set (Section A), replacing the “doing the sks gesture” with a natural language description of the class, e.g., for the class “Drumming Fingers” - “drumming finger”. We then classify the videos with the pre-trained network trained on the Jester dataset. No videos were correctly classified because (i) the gestures are not well represented in the model and (ii) the generated videos do not adhere to the spatial assumption of the Jester dataset (where a single person is performing a gesture in front of a camera).

### v2

@@ 4.2 Automated quantitative evaluation via recognition
| Real (WebVid10M) | 66.7 | 10.9 |
| Real (Jester) | 70.6 | 8.7 |
Strategy
| Tuning Method | Layers | Accuracy | Copy |
| LoRA | All | 17.4 | 0.5 |
| LoRA | Ours | 10.6 | 3.3 |
| Full | All | 68.7 | 17.0 |
| Full | Ours | 70.6 | 8.7 |
@@
First, we study the choice of the optimized parameters. As in the case of the image customization methods, it is clear that optimizing the spatial layers of the model is not enough to learn a novel motion pattern, as shown in Table 2(b). Training the key and value projections in the cross-attention layers of the spatial layers achieves a good balance in terms of accuracy and copying score, as shown Table 2(a). We observe a two-fold reduction in the copying score when training only those parameters and an 8% improvement in motion accuracy compared to not training any spatial parameters.
@@
Next, in Table 2(c), we consider the role of optimizing the text token embedding. We notice that when jointly training the text token embedding with the model’s parameters, the motion accuracy is higher than when training only the model parameters. However, it also leads to nearly three times as much memorization of the training video appearance, as evidenced by the significantly increased Copy score. We hypothesize that this indicates that the text token embedding is learning something related to the appearance of the training videos. To test our sampling strategy we compare a model trained with our motion pattern sampling strategy (“Coarse-noise”) to a model that simply samples diffusion time steps uniformly (“Uniform”). As shown in Table 2(d), our sampling strategy improves the motion accuracy and reduces the appearance copying. Following our intuition regarding learning motion patterns rather than fine-grained details, our model is less prone to replicating the training examples.
We study the effect of the regularization prior in Table 2(e). We train four models with the same design choice as our method, yet we choose the regularization set to be either: (i) empty, (ii) a synthetic set of videos generated from the original text-to-video model, or (iii) real videos from the WebVid10m dataset (iv) containing real videos that contain similar visual scenes. We observe a significant drop in performance both without using a prior as well as using a synthetic prior. Using real data is beneficial for fine-tuning. We observe that using motions and visual scenes similar to the target motion is optimal. However, should this support set be unavailable, off-the-shelf dataset is still better than no regularization or synthetic video. Overall, our method achieves a good balance between high motion accuracy and low copying score. Lastly, we compare full parameter fine-tuning to low-rank adaptation in Table 2(f). We test two variants; adding LoRA adapters to both spatial and temporal transformers (first row), and to the same layers as our NewMove (second row). In our experiments we use rank 32 and generate videos with LoRA scale 1, we observe that full parameter fine-tuning achieves higher motion accuracy. Qualitatively, we observe that LoRA models perform relatively well on coarser gestures like Drumming Fingers but struggle to learn fine-grained motions like Sliding Two Fingers Up. Quantitative evaluation of motion fidelity. We quantitatively evaluate our approach by computing metrics corresponding to the quality of the generated motion and the overall fidelity of the generated video with respect to the input text prompt. We compare to baseline methods that have explicitly learned the motion pattern from a few examples. We follow the evaluation protocol as defined in Section 4.2. We report the motion accuracy and text alignment metrics across different methods in Table 3 and show qualitative comparison (Figure 3).
We observe that Textual Inversion completely fails to learn the motion, this failure is potentially because the text encoder has been trained only on image-text pairs, and the embedding fails to generalize to the video domain. Additionally, the text embedding is injected only into spatial layers in the cross-attention layers, and not the temporal ones. Alternatively, because we are learning an unseen motion pattern, which is more complex than a static object, the embedding does not have enough parameters to learn it. Dreambooth and Custom Diffusion learn to adhere to the spatial structure of the training videos and some hand manipulation. However they fail to accurately reproduce the motion and produce a temporally smooth video. Our approach yields over a 2 improvement in motion accuracy and more faithfully adheres to the text prompt. We also compare with the concurrent MotionDirector method [53] and find that our approach outperforms it in motion accuracy and text alignment. However, with longer training, the baseline achieves similar performance (see Appendix for details and qualitative comparison).
@@ 5 Conclusion
We present an approach for motion customization in text-to-video diffusion models. Our method can learn a new motion pattern from a set of few exemplar videos of different subjects performing the same motion. We conduct a thorough ablation study that identifies the key components of the method and evaluate them with respect to motion accuracy. We demonstrate qualitative results of our method and evaluate our method with a user study, specifically in scenarios where our method generalizes the motion to unseen actors, multiple people performing the motion, and different viewpoints.